OpenStudy (anonymous):

Write the standard form of the equation of the conic section with the given charateristics. 1. Parabola with vertex at (0, 0) and directrix at x = 5. 2. Circle with center at (3, -1) and radius 4. 3. Ellipse with foci at (4, 2) and (10, 2) and co-vertices at (7, 0) and (7, 4) 4. Hyperbola with vertex at (0, 4), focus at (0, 5), and center at (0, 1)

OpenStudy (anonymous):

sketching parabola: Format: (y-k)^2 = 4p(x-h) I. Vertex is Pt(0,0) II. Opens left. Line of symmetry is the x-axis III. Directrix is x = 5 -p = 5 OR p = -5 and 4p = -20 y^2 = -20x OR x = (-1/20)y^2 The first question